# Learning Task Decomposition to Assist Humans in Competitive Programming

### v1

@@ Abstract
When using language models (LMs) to solve complex problems, humans might struggle to understand the LM-generated solutions and repair the flawed ones. To assist humans in repairing them, we propose to automatically decompose complex solutions into multiple simpler pieces that correspond to specific subtasks. We introduce a novel objective for learning task decomposition, termed assistive value (AssistV), which measures the feasibility and speed for humans to repair the decomposed solution. We collect a dataset of human repair experiences on different decomposed solutions. Utilizing the collected data as in-context examples, we then learn to critique, refine, and rank decomposed solutions to improve AssistV. We validate our method under competitive programming problems: under 177 hours of human study, our method enables non-experts to solve 33.3% more problems, speeds them up by 3.3x, and empowers them to match unassisted experts.
@@ 1 Introduction
With their increased capabilities, language models (LMs) are used to perform increasingly complex and high-impact problems (Trinh et al., 2024; Li et al., 2022; Huang et al., 2023b). The scalable oversight challenge emerges (Amodei et al., 2016): LMs might fail to provide reliable solutions for these problems, but it is also difficult for humans to evaluate and improve LMs’ solutions due to therequired significant time and expertise.
@@
However, not all decompositions are helpful, and it is challenging for humans to design an effective oneConnolly and Dean (1997); Selby (2015); Correa et al. (2023). For example, Charitsis et al. (2023) shows that improper decomposition designed by novice programmers can impede human debugging performance. To decompose better, we need methods beyond using fixed heuristic rules Wu et al. (2021) or learning from author-crafted demonstrations Yao et al. (2022); Zelikman et al. (2023).
In this paper, we introduce a novel objective for learning task decomposition: assistive value (AssistV, Equation 1), which measures the feasibility and speed of humans to repair a decomposed solution in the actual annotation process (Figure 1 right). To improve AssistV, we first collect a dataset of decompositions, measure their AssistV, and ask human annotators to provide a natural language critique on what makes a decomposition (not) helpful. Then we design a three-stage process to generate high-AssistV decompositions, where each stage is implemented by an LLM that learns from our dataset in context: 1) learn a critique model $\pi_{\text{critique}}$ for predicting human critique on how to improve the initial decomposition for higher AssistV, 2) learn a refine model $\pi_{\text{refine}}$ to incorporate the critique to refine the initial decomposition, and 3) learn a rank model $\pi_{\text{rank}}$ to select a decomposition with high AssistV.
@@ 5.5 Scalable Oversight
Advancements in LLMs have intensified the need for scalable, reliable human oversight on complex tasks that reach or even exceed the capabilities of human experts. Addressing this, prior research has explored various assistance methods, such as self-critiquing Saunders et al. (2022), AI debate Parrish et al. (2022), and decomposition Christiano et al. (2018); Wu et al. (2021). However, previous works overlook the feedback from actual assisted humans, potentially leading to reduced helpfulness of the assistance to humans Xie et al. (2016); Parrish et al. (2022). To bridge this gap, we propose to learn assistance models from human feedback.

### v2

@@ Abstract
When using language models (LMs) to solve complex problems, humans might struggle to understand the LM-generated solutions and repair the flawed ones. To better assist humans in repairing them, we propose to automatically decompose complex solutions into simpler pieces corresponding to specific subtasks. We introduce a novel objective for learning task decomposition, termed assistive value (AssistV), which measures the feasibility and speed for humans to repair the decomposed solution. We collect a dataset of human repair experiences on different decomposed solutions. Utilizing the collected data as in-context examples, we then learn to critique, refine, and rank decomposed solutions to improve AssistV. We validate our method under competitive programming problems: under 177 hours of human study, our method enables non-experts to solve 33.3% more problems, speeds them up by 3.3x, and empowers them to match unassisted experts.
@@ 1 Introduction
With their increased capabilities, language models (LMs) are used to perform increasingly complex and high-impact problems (Trinh et al., 2024; Li et al., 2022; Huang et al., 2023b). This however causes the scalable oversight challenge (Amodei et al., 2016): LMs might fail to provide reliable solutions for these problems, but it is also difficult for humans to evaluate and improve LMs’ solutions due to the required significant time and expertise.
@@
However, not all decompositions are helpful, and it is challenging for humans to design a generally effective one Connolly and Dean (1997); Selby (2015); Correa et al. (2023). For example, Charitsis et al. (2023) showed that improper decomposition designed by novice programmers can impede human debugging performance. To decompose better, we need methods beyond using fixed heuristic rules Wu et al. (2021) or learning from author-crafted demonstrations Yao et al. (2022); Zelikman et al. (2023).
In this paper, we introduce a novel objective for learning task decomposition: assistive value (AssistV, Eq (1)), which measures the feasibility and speed of humans to repair a decomposed solution in the actual annotation process (Figure 1 right). To improve AssistV, we first collect a dataset of decompositions, measure their AssistV, and ask human annotators to provide a natural language critique on what makes a decomposition (not) helpful. Then we design a three-stage process to generate high-AssistV decompositions, where each stage is implemented by an LLM that learns from our dataset in context: 1) learn a critique model $\pi_{\text{critique}}$ for predicting human critique on how to improve the initial decomposition for higher AssistV, 2) learn a refine model $\pi_{\text{refine}}$ to incorporate the critique to refine the initial decomposition, and 3) learn a rank model $\pi_{\text{rank}}$ to select a decomposition with high AssistV.
@@ 5.5 Scalable Oversight
Advancements in LLMs have intensified the need for scalable, reliable human oversight on complex tasks that reach or even exceed the capabilities of human experts. Addressing this, prior research has explored various assistance methods, such as self-critiquing Saunders et al. (2022), AI debate Parrish et al. (2022), using simple and informative examples zhong2023non, and decomposition Christiano et al. (2018); Wu et al. (2021). However, previous works overlook the feedback from actual assisted humans, potentially reducing helpfulness for humans Xie et al. (2016); Parrish et al. (2022). To bridge this gap, we propose to learn assistance models from human feedback.